OpenStudy (castiel):

Solve inequality |(x+3)^2|<1

OpenStudy (castiel):

The exponent is confusing me

OpenStudy (irishboy123):

get rid of the || first, its redundant then square root?!

OpenStudy (anonymous):

Let's solve your inequality step-by-step. (x+3)2<1 x2+6x+9<1 Let's find the critical points of the inequality. x2+6x+9=1 x2+6x+9−1=1−1(Subtract 1 from both sides) x2+6x+8=0 (x+2)(x+4)=0(Factor left side of equation) x+2=0 or x+4=0(Set factors equal to 0) x=−2 or x=−4 Check intervals in between critical points. (Test values in the intervals to see if they work.) x<−4(Doesn't work in original inequality) −4<x<−2(Works in original inequality) x>−2(Doesn't work in original inequality) Answer: −4<x<−2

OpenStudy (castiel):

so basically because of the exponent it's the same as |(x+3)|<1

OpenStudy (anonymous):

yeahh

OpenStudy (irishboy123):

no the square will be +ve so you don't need the | |, it's redundant so get rid of it first

OpenStudy (anonymous):

yeahh thats what throws it off my fault

OpenStudy (irishboy123):

Casteil try doing it yourself having got rid of the ||, we can addresss the exponent directly and in far far fewer lines!!!

OpenStudy (irishboy123):

without the ||'s, what does the equation now say?

OpenStudy (castiel):

but square root of (x+3)^3 is again |x+3|

OpenStudy (castiel):

^2 not ^3

OpenStudy (anonymous):

Right

OpenStudy (irishboy123):

if you know that \((x+3)^2 < 1\) then you know that \( |x+3|<1\), yes?

OpenStudy (castiel):

and then from that -1<x+3<1 and we get -4<x<-2

OpenStudy (anonymous):

yes

OpenStudy (irishboy123):

well done castiel!